OpenStudy (anonymous):

a lot measures 10 m by 26 m on its bases and 8m height. at the center of the lot is a house with floor dimensions of 12 by 10.5m. Find the area not covered by the house.

OpenStudy (shane_b):

What part are you having trouble with? You just need to calculate the area of the lot and subtract off the area of the house. The height in the questions is not used.

OpenStudy (anonymous):

@Shane_B so the answer is 134?

OpenStudy (shane_b):

Yes

OpenStudy (anonymous):

the unit is meters? or m^2?

OpenStudy (shane_b):

m^2

OpenStudy (anonymous):

why is the height not being used?

OpenStudy (shane_b):

Since you're subtracting m^2 from m^2, the answer must be in m^2 :)

OpenStudy (shane_b):

Because you don't need the height to calculate area....who cares how tall the house is? It takes up the same area on the lot.

OpenStudy (anonymous):

ah ok.

OpenStudy (aripotta):

it says the lot has height

OpenStudy (anonymous):

trapezoid HAHAAHA no?

OpenStudy (aripotta):

i believe it's another trapezoid lol. this doesn't make sense

OpenStudy (anonymous):

what do you think @sh

OpenStudy (shane_b):

Yea,..I don't know why it says the lot has height. The question ends up asking "Find the area not covered by the house." That can be calculated without any heights.

OpenStudy (aripotta):

if the lot is shaped like a trapezoid...

OpenStudy (anonymous):

yea it could be

OpenStudy (aripotta):

i'm sensing a theme here lol

OpenStudy (shane_b):

Ahh...you could be right. I was thinking 3D height...not a 2D trapezoid. At any rate, if it's a trapezoid calculate the area of that and subtract off the area of the house.

OpenStudy (anonymous):

my dream house is trapezoid shape :))

OpenStudy (shane_b):

hehe

OpenStudy (aripotta):

you got it from here, simon?

OpenStudy (anonymous):

i got 22 m^2

OpenStudy (shane_b):

I got 18m^2

OpenStudy (aripotta):

me too. 18 as well

OpenStudy (anonymous):

i will try again

OpenStudy (aripotta):

alright

OpenStudy (anonymous):

area of trapezoid = 184?

OpenStudy (shane_b):

\[\large A_{lot}=\frac{1}{2}(b_1+b_1)h=\frac{1}{2}(10m+26m)(8m)=144m^2\]

OpenStudy (anonymous):

guys! i'm so sorry. it is 20 not 10 x(((

OpenStudy (anonymous):

my mind is confused and i am hungry.

OpenStudy (anonymous):

sorry

OpenStudy (aripotta):

oh.

OpenStudy (shane_b):

\[\large A_{lot}=\frac{1}{2}(b_1+b_1)h=\frac{1}{2}(20m+26m)(8m)=184m^2\]

OpenStudy (anonymous):

and then the other one is 126

OpenStudy (shane_b):

Yep....so 184-126 m^2

OpenStudy (anonymous):

58!

OpenStudy (aripotta):

yup

OpenStudy (shane_b):

I think we have a winner ;)

OpenStudy (anonymous):

hayyy.. thanks you guys!

OpenStudy (aripotta):

no problem :D